INFRARED SPECTRA AND PHOTOISOMERIZATION OF HSO_{2} AND HOSO TRAPPED IN SOLID NEON

\underline{MARILYN E. JACOX} and WARREN E. THOMPSON, Optical Technology Division, National Institute of Standards and Technology, Gaithersburg, MD 20899-8441.

When a Ne:H_{2}:SO_{2} mixture is codeposited at 4.2 K with a beam of neon atoms that have been excited in a microwave discharge, the resulting solid deposit has infrared absorptions characteristic of several neutral and ionic products. Among these are HSO_{2} and cis-HOSO, recently identified among the products of the reaction of photoproduced H atoms with SO_{2} in the heavier rare gas matrices ^{a} . The assignments and the photoisomerization behavior reported in the earlier study are supported by the neon-matrix observations, which include studies of the sulfur-34 and oxygen-18 substituted species. After photodestruction of the HSO_{2} and cis-HOSO, slight warmup of the deposit to induce H-atom migration leads to regeneration of HSO_{2}.

aE. Isoniemi, L. Khriachtchev, J. Lundell, and M. Räsänen, Phys. Chem. Chem. Phys. 4, 1549 (2002).